# The Multi-Range Theory of Translation Quality Measurement: MQM scoring models and Statistical Quality Control

### v3

@@ 1 Introduction and Background
Translation errors often carry risks. Their consequences range from minor misunderstandings to serious legal, financial, reputational, or health-related outcomes for end users, translation providers, clients, and other stakeholders. Risk mitigation requires evaluation to identify and quantify these risks. The Multidimensional Quality Metrics (MQM) framework for analytic Translation Quality Evaluation (TQE) was first proposed by Lommel et al., 2014a just before the arrival of NMT, originally published as a deliverable of the EU-funded QTLaunchpad project. From the very beginning, it was designed for evaluation of both human translation (HT), and machine translation (MT), and it can now be applied to AI-generated translation. MQM has formalized and standardized the so-called analytic approach to translation quality measurement.
@@
Nevertheless, the MQM framework has been picked up again by the leading MT shared task venue WMT since 2021 as the official human evaluation strategy to judge the submitted MT systems (Freitag et al.,, 2021, 2022, 2023). There have been automatic evaluation designs that aim to mimic the MQM idea, such as COMET-MQM reported from WMT2020 metrics shared task Mathur et al., (2020).
@@
In this paper, we introduce the latest developments of the MQM framework from the MQM Council which comprises a voluntary, community-driven research and standardization group composed of experts interested in translation quality evaluation, who have been developing the MQM since 2016.  [^1] https://themqm.org/mqm-council/
@@
In recent years, the widely used DQF subset of MQM has been improved and updated to become MQM Core. This error typology is better adapted to quality management systems with a clearer structure for devising improvement actions. The latest iteration of the framework includes the revised MQM Core and MQM Full error typology, a new linear scoring model with calibration, a process description with a sample scorecard, and now also a non-linear MQM scoring model. (appendix 7).
@@
We further argue that different evaluation approaches have to be used for three ranges of sample sizes. For the first time in the history of the translation and localization industry, to the best of our knowledge, this paper presents a multi-range, versatile theory and technique of Translation Quality Evaluation, making it possible for interested researchers to construct almost any analytic metrics derived from this approach. We also explain in this paper why segment-level scores cannot be accurate in principle, and explain the area of applicability of Statistical Quality Control to Translation Quality Evaluation.
2 Sampling
@@
For any statistical approach to be applicable, it is critical to know what statistical distribution is valid before choosing the right distribution for further statistical analysis. Such analysis is based on the notion that (a) all errors are independent and (b) the probability of errors in the text is uniform. Although these assumptions are always made, both assumptions are in fact not true with respect to texts (and translation products) in general.
@@
Practitioners in the language and translation industry know that translation errors are not, in general, uniformly distributed in content and, what is more, over time. Furthermore, their significance and “weights” are also different in various parts of the material and/or types of material, and also vary according to other sometimes unpredictable factors. Furthermore, different types of errors may depend on each other. Indeed, some errors only occur when triggered by other fundamental errors.
@@
There are about a dozen very important factors which can influence why error distribution in text is not, in fact, uniform. Because of this, it is always recommended to revise, review, and evaluate the entire text. But no one has the resources to fully evaluate everything. That’s why translation quality evaluation is typically based on evaluating samples. Sometimes this sampling is done by selecting full-text samples from a larger population, e.g. evaluating every tenth translation fully, or selecting complete shorter texts to arrive at the determined sample size. A typical approach in a traditional localization process is to select samples of medium sizes (500-5000 words) and then apply an analytical approach. However, smaller samples and larger samples are also possible, and it is important to consider the entire range of possible sample sizes, from one sentence to very large documents, which is the purpose of this paper.
@@ 3 Low IRR is not a bug, but a feature
Human translation quality evaluations on small samples have low Inter-Rater Reliability (IRR). What matters in human evaluation is that trained linguists tend to agree (demonstrate high IRR) on bigger samples, but on a segment level two linguists are bound to have three opinions on the same issue. This phenomenon has been widely recognized by both experts in the translation industry and data science. Research in the field of translation error annotation has demonstrated that linguists disagree even on the precise scope of errors, as well as about error categorizations and severity attributions (Lommel et al., 2014b ). The researchers established that these issues are not a consequence of low linguist qualifications or technical problems. They are inherent in language itself.
@@
Text is not the information itself; it is merely the conveyor of the expression of intended meaning. Moreover, not only can we not know for sure the author’s intended meaning, which leaves the text open to the reader’s interpretation. This interpretation depends on the reader’s cultural, educational, and professional background and experiences, as well as on their skills and the context of the communication act. Additionally, language is highly ambiguous, offering numerous expression tools that allow a single sentence to be interpreted in many ways. This inherent ambiguity leads to significant uncertainty in any error annotation, which is a fundamental property of language rather than a flaw of human assessors.
@@
Data scientists often make shallow conclusions from low IRR, believing that automatic calculation of errors will resolve this problem. In recent years, many automatic metrics have been constructed, often claiming “human judgment” quality. Lately, AI-based metrics have appeared, along with unverified and unproven claims that “GenAI seems to be capable of measuring translation quality”. In reality, however, GenAI does not ”think” and does not ”understand” anything and for this reason, the factual accuracy of GenAI generation remains a huge problem. No language model is Turing complete. Language models are constrained by their architecture and the limitations of their training data and computational resources. Language models are specialized tools designed for processing and generating text based on patterns learned from vast amounts of data. They are good at some tasks such as language generation, and summarization, but they are not capable of performing arbitrary computations or reasoning in the same way that a Turing complete system can. Because of this, language models are not considered general-purpose reasoners. They can provide useful responses and assist with many tasks within their scope of capabilities, but they cannot reason and compute in the general-purpose manner that a Turing machine or Turing complete system can. Among other things, the factual accuracy of GenAI generation remains to be a huge problem. This affects both error annotation capability and accuracy on a secondary process level. GenAI may produce fluent text, but it performs worse on derivative tasks, such as error annotation tasks. In other words, GenAI misses factual errors when generating content, and of course, does not “see” errors during the annotation process, which also results in issues with the accuracy of this process. Additionally, the better the GenAI output is, the more variable it becomes, which is similar (although different in nature) to the variability in human judgment. Human judgment is variable because different people may have their own interpretation of the text. Advanced GenAI behaves similarly, with its variable output becoming another interpretation. However, this interpretation is not verified or supported by human intelligence.
@@
For all we know, various automatic and GenAI measurement results must be validated by analytic human evaluation on samples of sufficient sizes, which converges to higher IRR (inter-rater reliability) in controlled settings with training and monitoring.
@@ 4 Three sample-size ranges – three very different methods
The industrial era has developed a vast mathematical and methodological apparatus for measuring the quality of products by evaluating small samples from very large production lots. In a setting where decisions about the quality of a large lot are based on small samples, the uncertainty is so high that only sophisticated methods of statistical quality control can handle such a problem. They have been extensively developed and described by many researchers, such as Montgomery, (2019), and have been long standardized by ISO  [^2] ISO 2859-2:2020 Sampling procedures for inspection by attributes https://www.iso.org/standard/64505.html.
In a nutshell, errors always have a statistical nature, and this is yet another reason why segment-level quality scores do not make sense – not only the methods of producing them must be properly benchmarked and verified, but the fundamental uncertainty of individual error annotation is so high that it is not possible to get a ”true” measurement except for very trivial mechanical spelling errors.
@@
When measuring errors on a very small sample, you need to apply the Statistical Quality Control (SQC) method. These methods are extremely complex due to the underlying statistical and mathematical apparatus, which is why ISO 2859-1 consists of 94 pages of tables. Additionally, the conclusions drawn from applying these methods are not quality ratings but rather probabilities of the producer’s and consumer’s risks, which can be determined using such methods.
@@
In an earlier paper “Measuring the uncertainty in Translation Quality Evaluation” (Gladkoff et al., (2022)), it was demonstrated that for the translation quality rating to have low variability, the sample size should be greater than 200 sentences (3400 words). As you go to smaller sample sizes, the confidence interval explodes. This means that even a 2000-word sample introduces significant uncertainty in the quality measurement result, and even more so with samples of 1000 words and especially 500 words.
@@
The lowest known credible metric is the ATA certification model, where linguists are asked to translate one page (250 words). To address the uncertainty of evaluating such a small sample, two different reviewers assess the work, effectively duplicating the evaluation effort to mitigate the risks associated with the limited sample size. Additionally, the ATA certification requires linguists to translate two one-page samples to see how they handle the translation of different types of text, which also helps reduce the uncertainty of evaluating small samples.
@@
The methods of Statistical Quality Control are outside the scope of this paper. However, for this article, it is important to note that the translation quality of a sample smaller than 15-17 sentences (one page, 250 words) falls into the realm of SQC and cannot be measured by analytic quality evaluation methods unless the sample covers the entire text.
@@
For samples of approximately 300 words and above, the effects of statistical uncertainty and low Inter-Rater Reliability cease to have a significant adverse effect, making methods of analytical quality evaluation applicable. For samples larger than approximately 5000 words, other effects start to manifest, such as the priming effect of human perception. Non-linearity starts to become apparent with larger samples. In this paper, we introduce a non-linear calibration model that works for samples of both medium (customary) and large sizes. The three ranges of sample sizes are governed by entirely different mathematical apparatus. These ranges are shown in Figure 8.
@@ 5.2 Scoring Model
The Scoring Model with Calibration works with all the same original values and parameters as the Raw Scoring Model, but a few additional parameters are required as explained in Score Calculation with Calibration (6.6).
@@ 5.3 Setting up an MQM evaluation system
To set up the MQM evaluation process, implementers must first create a specific metric by:
  • Evaluating translation project specifications and analyzing them concerning end-user needs;
  • Selecting Error Types from the MQM Error Typology that are appropriate to the defined specifications to create the error typology for the metric;
  • Integrating the created MQM metric, suitable for their specific use case, into a scoring system into their tool environment or using it in a scorecard format;
  • Determining sample size ranges and defining sampling procedures.
Historically, translation evaluators have used spreadsheet-like scorecards (see Figure 1 below), but today evaluators are more likely to annotate errors and integrate scoring calculations in specialized tools or utilities built into their translation environment. To visualize and describe the annotation and score-generation process, it is helpful to use a scorecard as an example. The sample scorecards used here represent useful options, but they are not normative.
The default setting for MQM is to use four severity levels, with the severity multipliers 0-1-5-25. Implementers can choose to use different ranges of severity levels, as explained later (6.3.3), and implement them via their respective Severity Penalty Multiplier.
6 Parameters of the Evaluation Quality Metric
6.1 Calculation Values
The sample scorecard in Figure 1 shows the results of applying one possible metric to a specific evaluation task for a sample of segments (the size of which is stated in the Evaluation Word Count). Rows 29–35 list the selected Error Types – in this case the seven high-level error dimensions of MQM Core.
The following values highlighted in Figure 1 play major roles in designing translation quality evaluation models. The abbreviations listed below are sometimes used when discussing formal equations.
6.1.1 Evaluation Word Count (EWC)
6.1.2 Reference Word Count (RWC)
6.2 Maximum Score Value (MSV)
6.2.1 Passing Threshold (PT)
6.2.2 Defined Passing Interval (DPI)
6.2.3 Final Quality Rating (Pass/Fail)
6.2.4 Error Type Weight (ETW)
6.3 Error Annotation Values
6.3.1 Error Type Number (ET No)
6.3.2 Error Severity Level
6.3.3 Severity Penalty Multiplier
6.4 Scoring Models Parameters
6.4.1 Error Count (EC)
6.4.2 Error Type Penalty Total (ETPT)
6.4.3 Absolute Penalty Total (APT)
6.4.4 Per-Word Penalty Total (PWPT)
6.4.5 Normed Penalty Total (NPT)
6.4.6 Quality Score (QS)
6.5 Calculating the Linear Quality Score
6.5.1 Quality Score without Calibration (Raw Score)
6.5.2 Quality Score with Calibration
@@
The process of calibration acts like a magnifying glass for viewing the otherwise very small or inconsistent ranges of acceptance close to 100. This approach makes the quality rating easier to use and understand, highlighting differences in translation quality for evaluated texts more clearly.
6.6 Score Calculations
6.6.1 Calculating the raw quality scores
Calculating scores without calibration uses the steps shown on Figure 3. For a list of all parameters and their abbreviations, see the Appendix: Scoring Model Parameters.
@@ 6.6.1 Calculating the raw quality scores
6.6.2 Calculating the quality scores with calibration
The scoring formula for calculating the quality score with calibration works with the standard calculation values, such as Evaluation Word Count, Absolute Penalty Total and Normed Penalty Total. However, a few additional values and parameters have to be defined. These are used to pre-define the specified acceptance criteria (the error tolerance) and to link these criteria to a scale that should be understandable or appropriate for all stakeholders.
The following values used in the score calculation above are pivotal for a score calibrated with respect to a predefined Passing Threshold.
Scaling Factor (SF) Parameter to scale the Acceptable Penalty Points (APP) for the reference word count across the Defined Passing Interval (DPI). Let’s consider Figure 2. On the top Raw Scale the raw passing threshold is 98, which means that a maximum of 20 raw penalty points are allowed on a sample of 1000 words. On a Calibrated Score scale (bottom) the Defined Passing Interval (DPI) is 15 (upward from 85 to 100). Therefore, the raw Passing Interval scales down from 20 on the raw scale to 15 on the Calibrated Scale. The Scaling Parameter for this new value will be 15/20=0.75. On the Raw Scale, the raw score is 98.44, which means that NPT=15.6 (the error density of a sample is equivalent to 15.6 errors on 1000 words). The trick to the Calibration is that what is scaled is not a passing interval, but rather, NPT. We multiply raw NPT of 15.6 by a scaling factor 15/20=0.75, which resolves in NPT = 11.7 on the calibrated scale. Therefore, the calibrated score will be 100-11.7=88.3 (bottom scale).
7 Non-Linearity
This section describes the non-linear scoring model which solves the problem of the non-linearity of human perception with respect to samples of medium to large size.
All is well with linear scoring until the sample size is either too small (requiring methods of Statistical Quality Control) or too large. Many practitioners have observed that linear scoring defined for small to medium samples does not work for larger chunks of text.
The reason that changing human perception fails a linear model is that the human mind is not a fixed state machine – we change our behaviour and our opinions with experience. This effect is known in cognitive science as “priming,” where encountering certain features in the text makes readers more alert to them as they work their way through the text.
Priming works on errors too. When we see two errors on one page, then two errors on the next page, and three errors on the page thereafter, we get the impression that there are simply too many errors in the document.
This is observed in real translation quality evaluation practice as follows (this is an actual quote from a very large buyer of translation services):
”Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them. We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient. The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula. For example, we might feel that if a translation sample is about one page, a single major mistranslation error is enough to say that it fails. However, if the sample is seven pages, we are not okay with allowing seven major mistranslation errors before it fails. Instead, we would prefer to fail the sample already at three or four errors.
This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.”

### v4

@@ 1 Introduction and Background
Translation errors often carry risks. Their consequences range from minor misunderstandings to serious legal, financial, reputational, or health-related outcomes for end users, translation providers, clients, and other stakeholders (Han et al.,, 2024). Risk mitigation requires evaluation to identify and quantify these risks. The Multidimensional Quality Metrics (MQM) framework for analytic Translation Quality Evaluation (TQE) was first proposed by Lommel et al., 2014a just before the arrival of NMT, originally published as a deliverable of the EU-funded QTLaunchpad project. From the very beginning, it was designed for evaluation of both human translation (HT), and machine translation (MT), and it can now be applied to AI-generated translation. MQM has formalized and standardized the so-called analytic approach to translation quality measurement.
@@
Nevertheless, the MQM framework has been picked up again by the leading MT shared task venue WMT since 2021 as the official human evaluation strategy to judge the submitted MT systems (Freitag et al.,, 2021, 2022, 2023). There have been automatic evaluation designs that aim to mimic the MQM idea, such as COMET-MQM reported from WMT2020 metrics shared task (Mathur et al.,, 2020).
@@
In this paper, we introduce the latest developments of the MQM framework from the MQM Council which comprises a voluntary, community-driven research and standardization group composed of experts interested in translation quality evaluation, who have been developing the MQM since 2016.  [^1] https://themqm.org/mqm-council/ These start with the discussion of Sampling and Low IRR phenomena, followed with sample-sizes, MQM2, Projecting PI and Formulas, and the introducing non-linearity. We leave the detailed MQM Parameters into appendix for indexing.
@@
In recent years, the widely used DQF subset of MQM has been improved and updated to become MQM Core. This error typology is better adapted to quality management systems with a clearer structure for devising improvement actions. The latest iteration of the framework includes the revised MQM Core and MQM Full error typology, a new linear scoring model with calibration, a process description with a sample scorecard, and now a non-linear MQM scoring model. (appendix 7).
@@
We further argue that different evaluation approaches have to be used for three ranges of sample sizes. For the first time in the history of the translation and localization industry, to the best of our knowledge, this paper presents a multi-range, versatile theory and technique of Translation Quality Evaluation, making it possible for interested researchers to construct almost any analytic metrics derived from this approach. We also explained in this paper why segment-level scores cannot be accurate in principle, and explain the area of applicability of Statistical Quality Control to Translation Quality Evaluation.
2 On Sampling
@@
For any statistical approach to be applicable, it is critical to know what statistical distribution is valid before choosing the right distribution for further statistical analysis. Such analysis is based on the notion that (a) all errors are independent and (b) the probability of errors in the text is uniform. Although these assumptions are always made, both assumptions are not true concerning texts (and translation products) in general.
@@
Practitioners in the language and translation industry know that translation errors are not, in general, uniformly distributed in content and, what is more, over time. Furthermore, their significance and “weights” are also different in various parts of the material and/or types of material, and also vary according to other sometimes unpredictable factors. In addition, different types of errors may depend on each other. Indeed, some errors only occur when triggered by other fundamental errors.
@@
There are about a dozen very important factors that can influence why error distribution in text is not, in fact, uniform. Because of this, it is always recommended to revise, review, and evaluate the entire text. But no one has the resources to fully evaluate everything. That’s why translation quality evaluation is typically based on evaluating samples (Gladkoff et al.,, 2022). Sometimes this sampling is done by selecting full-text samples from a larger population, e.g. evaluating every tenth translation fully, or selecting complete shorter texts to arrive at the determined sample size. A typical approach in a traditional localization process is to select samples of medium sizes (500-5000 words) and then apply an analytical approach. However, smaller samples and larger samples are also possible, and it is important to consider the entire range of possible sample sizes, from one sentence to very large documents, which is the purpose of this paper.
@@ 3 Low IRR is not a bug, but a feature
Human translation quality evaluations on small samples have low Inter-Rater Reliability (IRR). What matters in human evaluation is that trained linguists tend to agree (demonstrate high IRR) on bigger samples, but on a segment level two linguists are bound to have three opinions on the same issue. This phenomenon has been widely recognized by both experts in the translation industry and data science (Gladkoff et al., 2023b, ). Research in the field of translation error annotation has demonstrated that linguists disagree even on the precise scope of errors, as well as about error categorizations and severity attributions (Lommel et al., 2014b, ). The researchers established that these issues are not a consequence of low linguist qualifications or technical problems. They are inherent in language itself.
@@
Text is not the information itself; it is merely the conveyor of the expression of intended meaning. Moreover, not only can we not know for sure the author’s intended meaning, which leaves the text open to the reader’s interpretation. This interpretation depends on the reader’s cultural, educational, and professional background and experiences, as well as on their skills and the context of the communication act. Additionally, language is highly ambiguous, offering numerous expression tools that allow a single sentence to be interpreted in many ways. This inherent ambiguity leads to significant uncertainty in any error annotation, which is a fundamental property of language rather than a flaw of human assessors (Gladkoff et al.,, 2022).
@@
Data scientists often make shallow conclusions from low IRR, believing that automatic calculation of errors will resolve this problem. In recent years, many automatic metrics have been constructed, often claiming “human judgment” quality. Recently, AI-based metrics have appeared, along with unverified and unproven claims that “GenAI seems capable of measuring translation quality” (Gladkoff et al., 2023a, ). In reality, however, GenAI does neither “think” nor “understand” anything and for this reason, the factual accuracy of GenAI generation remains a huge problem. No language model is Turing complete  [^2] https://ai-lab.logrusglobal.com/why-no-agi-can-be-built-with-language-models/. Language models are constrained by their architecture and the limitations of their training data and computational resources. Language models are specialized tools designed for processing and generating text based on patterns learned from vast amounts of data. They are good at some tasks such as language generation, and summarization, but they are not capable of performing arbitrary computations or reasoning in the same way that a Turing complete system can. Because of this, language models are not considered general-purpose reasoners. They can provide useful responses and assist with many tasks within their scope of capabilities, but they cannot reason and compute in the general-purpose manner that a Turing machine or Turing complete system can. Among other things, the factual accuracy of GenAI generation remains to be a huge problem. This affects both error annotation capability and accuracy on a secondary process level. GenAI may produce fluent text, but it performs worse on derivative tasks, such as error annotation tasks. In other words, GenAI misses factual errors when generating content, and of course, does not “see” errors during the annotation process, which also results in issues with the accuracy of this process. Additionally, the better the GenAI output is, the more variable it becomes, which is similar (although different in nature) to the variability in human judgment. Human judgment is variable because different people may have their own interpretation of the text. Advanced GenAI behaves similarly, with its variable output becoming another interpretation. However, this interpretation is not verified or supported by human intelligence.
@@
For all we know, various automatic and GenAI measurement results must be validated by analytic human evaluation on samples of sufficient sizes, which converges to higher IRR (inter-rater reliability) in controlled settings with training and monitoring (Gladkoff et al.,, 2022; Han et al.,, 2024).
@@ 4 Three sample-size ranges – three very different methods
The industrial era has developed a vast mathematical and methodological apparatus for measuring the quality of products by evaluating small samples from very large production lots, e.g. using student’s t-distribution by Gladkoff et al., 2023b . In a setting where decisions about the quality of a large lot are based on small samples, the uncertainty is so high that only sophisticated methods of statistical quality control can handle such a problem. They have been extensively developed and described by many researchers, such as Montgomery, (2019), and have been long standardized by ISO  [^3] ISO 2859-2:2020 Sampling procedures for inspection by attributes https://www.iso.org/standard/64505.html. In a nutshell, errors always have a statistical nature, and this is yet another reason why segment-level quality scores do not make sense – not only the methods of producing them must be properly benchmarked and verified, but the fundamental uncertainty of individual error annotation is so high that it is not possible to get a “true” measurement except for very trivial mechanical spelling errors.
@@
When measuring errors on a very small sample, we need to apply the Statistical Quality Control (SQC) method. These methods are extremely complex due to the underlying statistical and mathematical apparatus, which is why ISO 2859-1 consists of 94 pages of tables. Additionally, the conclusions drawn from applying these methods are not quality ratings but rather probabilities of the producer’s and consumer’s risks, which can be determined using such methods.
@@
In an earlier work by Gladkoff et al., (2022), it was demonstrated that for the translation quality rating to have low variability, the sample size should be greater than 200 sentences (3400 words). The authors discuss that if they go to smaller sample sizes, the confidence interval explodes. This means that even a 2000-word sample introduces significant uncertainty in the quality measurement result, and even more so with samples of 1000 words and especially 500 words.
@@
The lowest known credible metric is the ATA certification model (Han and Gladkoff,, 2022), where linguists are asked to translate one page (250 words). To address the uncertainty of evaluating such a small sample, two different reviewers assess the work, effectively duplicating the evaluation effort to mitigate the risks associated with the limited sample size. Additionally, the ATA certification requires linguists to translate two one-page samples to see how they handle the translation of different types of text, which also helps reduce the uncertainty of evaluating small samples.
@@
The methods of Statistical Quality Control (SQC) are outside the scope of this paper. However, for this article, it is important to note that the translation quality of a sample smaller than 15-17 sentences (one page, 250 words) falls into the realm of SQC and cannot be measured by analytic quality evaluation methods unless the sample covers the entire text.
@@
For samples of approximately 300 words and above, the effects of statistical uncertainty and low Inter-Rater Reliability cease to have a significant adverse effect, making methods of analytical quality evaluation applicable. For samples larger than approximately 5000 words, other effects start to manifest, such as the priming effect of human perception. Non-linearity starts to become apparent with larger samples. In this paper, we introduce a non-linear calibration model that works for samples of both medium (customary) and large sizes. The three ranges of sample sizes are governed by entirely different mathematical apparatus. These ranges are shown in Figure 7 (word counts: 0, 500, 5000+).
@@ 5.2 Scoring Model
The Scoring Model with Calibration works with all the same original values and parameters as the Raw Scoring Model, but a few additional parameters are required as explained in Score Calculation with Calibration (A.6).
@@ 5.3 Setting up an MQM evaluation system
To set up the MQM evaluation process, implementers must first create a specific metric by: 1) Evaluating translation project specifications and analyzing them concerning end-user needs; 2) Selecting Error Types from the MQM Error Typology that are appropriate to the defined specifications to create the error typology for the metric; 3) Integrating the created MQM metric, suitable for their specific use case, into a scoring system into their tool environment or using it in a scorecard format; 4) Determining sample size ranges and defining sampling procedures.
Historically, translation evaluators have used spreadsheet-like scorecards (see Figure 9), but today evaluators are more likely to annotate errors and integrate scoring calculations in specialized tools or utilities built into their translation environment. To visualize and describe the annotation and score-generation process, it is helpful to use a scorecard as an example. The sample scorecards used here represent useful options, but they are not normative.
The default setting for MQM is to use four severity levels, with the severity multipliers 0-1-5-25. Implementers can choose to use different ranges of severity levels, as explained later (A.3.3), and implement them via their respective Severity Penalty Multiplier. (Appendix A for detailed parameters)
6 Projecting PI and the Formulas
Figure 1 displays the projecting of the small passing interval window in the raw-score score to the scale of the calibrated score. Figure 2 introduces the formulas for raw score calculation.
@@ 6 Projecting PI and the Formulas
7 On Non-Linearity
This section describes the non-linear scoring model which solves the problem of the non-linearity of human perception concerning samples of medium to large size.
All is well with linear scoring until the sample size is either too small (requiring methods of Statistical Quality Control) or too large. Many practitioners have observed that linear scoring defined for small to medium samples does not work for larger chunks of text. The reason that changing human perception fails a linear model is that the human mind is not a fixed state machine – we change our behaviour and our opinions with experience. This effect is known in cognitive science as “priming,” where encountering certain features in the text makes readers more alert to them as they work their way through the text.
Priming works on errors too. When we see two errors on one page, then two errors on the next page, and three errors on the page thereafter, we get the impression that there are simply too many errors in the document. This is observed in real translation quality evaluation practice as follows (this is an actual quote from a very large buyer of translation services): Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them. We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient. The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula. For example, we might feel that if a translation sample is about one page, a single major mistranslation error is enough to say that it fails. However, if the sample is seven pages, we are not okay with allowing seven major mistranslation errors before it fails. Instead, we would prefer to fail the sample already at three or four errors. This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.”